### v1

# Calling a Spade a Heart: Gaslighting Multimodal Large Language Models via Negation

Abstract
Multimodal Large Language Models (MLLMs) have exhibited remarkable advancements in integrating different modalities, excelling in complex understanding and generation tasks. Despite their success, MLLMs remain vulnerable to conversational adversarial inputs, particularly negation arguments. This paper systematically evaluates state-of-the-art MLLMs across diverse benchmarks, revealing significant performance drops when negation arguments are introduced to initially correct responses. We show critical vulnerabilities in the reasoning and alignment mechanisms of these models. Proprietary models such as GPT-4o and Claude-3.5-Sonnet demonstrate better resilience compared to open-source counterparts like Qwen2-VL and LLaVA. However, all evaluated MLLMs struggle to maintain logical consistency under negation arguments during conversation. This paper aims to offer valuable insights for improving the robustness of MLLMs against adversarial inputs, contributing to the development of more reliable and trustworthy multimodal AI systems.
Figure 1: Examples demonstrate that GPT-4o initially provides correct answers but incorrectly revises its responses when confronted with user-provided negation arguments. GPT-4o shows a tendency to accept misleading inputs, often generating hallucinated explanations to justify the revised answers, a behavior that can be described as a form of “gaslighting".
1 Introduction
Multimodal Large Language Models (MLLMs) have achieved remarkable progress in understanding and generating language grounded in multimodal contexts, such as visual and textual inputs yin2023survey ; liu2024visual ; bai2023qwen ; hurst2024gpt . These models leverage cutting-edge advancements in Large Language Models (LLM) zhao2023survey and multimodal learning radford2021learning ; li2023blip ; liu2024improved , enabling them to excel in diverse tasks, such as image understanding, visual question answering, and multimodal reasoning. Recent breakthroughs, such as the introduction of proprietary models GPT-4o hurst2024gpt and Claude-3.5-Sonnet anthropic2024claude , have demonstrated exceptional performance, establishing new standards in multimodal understanding and reasoning. Additionally, open-source models like DeepSeek liu2024deepseek are closing the performance gap with the proprietary models, reflecting the rapid and competitive evolution of MLLMs. These advancements not only highlight the growing capabilities of MLLMs, but also suggest that the development trajectory is edging closer to achieving Artificial General Intelligence (AGI).
Despite the impressive advancements, MLLMs exhibit significant vulnerabilities when navigating complex conversational challenges, particularly those involving adversarial negation. This issue becomes evident when models struggle to critically analyze and resist unfaithful arguments, resulting in erroneous reversals of their initially correct answers wang2023can ; zhao2025aligning . As illustrated in Figure 1, GPT-4o initially provides correct answers for all examples. However, when presented with negation arguments by the user, GPT-4o often accepts these arguments, revising its answers incorrectly. Moreover, it even generates hallucinated explanations to justify these erroneous responses. Such susceptibility not only undermines their reliability but also exposes fundamental weaknesses in their reasoning and alignment mechanisms. These limitations are particularly concerning in high-stakes applications like healthcare diagnostics, autonomous decision-making, and content moderation, where the ability to maintain logical consistency and resist manipulation is crucial. Addressing these challenges requires a deeper understanding of how MLLMs process and align multimodal inputs, paving the way for more reliable and trustworthy models huang2024trustllm ; liu2023trustworthy .
Figure 2: Performance comparison of selected Multimodal Large Language Models (MLLMs) across eight benchmarks-MME fu2023mme , MMMU yue2023mmmu , MMMUPro yue2024mmmu-MMMUPro , MMBench liu2025mmbench , PoPE li2023evaluating-Pope , ChartQA masry2022chartqa , AI2Diagram kembhavi2016diagram and MathVista lu2023mathvista )-showing average accuracy before (initial answers) and after the introduction of negation arguments. Note that GPT-4o (accessed from 20-Jan-2025 to 22-Jan-2025) and Claude-3.5-Sonnet (accessed from 20-Jan-2025 to 23-Jan-2025) results are averaged over six benchmarks, excluding POPE and MMBench. Results for Gemini-1.5-flash are based on access from 13-Jan-2025 to 18-Jan-2025. More detail results can be found in Table 1.
To comprehensively understand the limitations of MLLMs in handling negation, we conduct extensive evaluations across multiple major benchmarks and state-of-the-art models. The benchmarks cover diverse datasets, ranging from general multimodal datasets such as MMMU yue2023mmmu and MMBench liu2025mmbench , chart dataset such as ChartQA masry2022chartqa , to math dataset MathVista lu2023mathvista , providing a robust framework for assessing the performance of MLLMs under negation arguments in conversation. The selected proprietary models, such as GPT-4o and Claude-3.5-Sonnet, alongside open-source counterparts like Qwen2-VL and LLaVA, allowed for a comparative analysis of reliability and accuracy. Through the extensive evaluation, we highlight the widespread vulnerability of MLLMs to negation arguments, with varying degrees of impact observed across different datasets and MLLMs.
Our observations reveal critical insights into the behavior of MLLMs. Notably, while larger models such as Qwen2-VL-72B-Instruct demonstrate higher accuracy in initial responses before negation, it also exhibit significant performance drops after negation, suggesting potential overfitting or alignment issues during training. On the contrary, proprietary models such as Gemini-1.5-flash, show better robustness compared to open-source models, but even they struggle to consistently defend correct answers against misleading negation arguments. These findings emphasize the urgent need for enhanced adversarial training techniques and robust alignment mechanisms to improve the reliability and integrity of MLLMs in real-world applications.
2 Evaluating Multimodal Large Language Model via Negation
2.1 Experimental Setup
Multimodal Large Language Models. To establish a comprehensive evaluation of the negation challenges posed to Multimodal Large Language Models (MLLMs), we assess a range of state-of-the-art models representing both proprietary and open-source systems. These models were chosen to reflect a diverse set of training approaches and capabilities in the multimodal AI landscape.
Proprietary Models: Our evaluation included leading proprietary models such as GPT-4o hurst2024gpt , Claude-3.5-Sonnet anthropic2024claude , and Gemini-1.5-flash team2024gemini1.5 . These models represent the most powerful multimodal AI capabilities.
Open-source Models: In addition to proprietary models, we included open-source models such as LLaVA-1.6-7B liu2024llavanext-llava1.6 , Qwen2-VL-7B-Instruct Qwen2VL-7B , LLaVA-NeXT-8B li2024llavanext-strong and Qwen2-VL-72B-Instruct Qwen2VL-7B . These open-source models showcase the rapid progress in publicly available multimodal AI systems.
Benchmarks. To comprehensively assess the robustness of MLLMs in handling negation, we employ a diverse set of benchmark datasets, including general multimodal datasets MME fu2023mme , MMMU yue2023mmmu , MMMUPro yue2024mmmu-MMMUPro , MMBench liu2025mmbench and PoPE li2023evaluating-Pope , Chart dataset ChartQA masry2022chartqa , diagram dataset AI2Diagram kembhavi2016diagram as well as Math dataset MathVista lu2023mathvista .
Evaluation Pipeline. To evaluate the robustness of MLLMs, we employ a structured evaluation pipeline designed to measure the impact of negation on model performance. The process consists of three key steps: First, questions and corresponding options (if applicable) from the original datasets are provided to the MLLMs, and their initial responses are recorded. Second, if the model’s initial answer aligns with the correct answer from the dataset, we introduce a negation argument challenging the validity of the original answer. The model’s response to the negation argument is recorded to assess whether it revises its original correct answer. If the model’s initial answer is incorrect, no negation argument is introduced, and the response remains unchanged. Third, we compute the accuracy of the model both before and after the introduction of negation arguments to quantify the drop in performance caused by adversarial negation. This evaluation pipeline enables a detailed analysis of how well MLLMs can maintain consistent reasoning under adversarial conversational challenges, shedding light on their robustness and potential vulnerabilities.
Negation Argument Generation. We systematically generate negation arguments tailored to different question formats using Llama3-8B-Instruct: (1) Yes/No Questions. Negation arguments are generated by rewriting the original question into an opposite statement based on the answer. If the answer is “Yes.", the question is rewritten as a negative statement and vice versa. For example, given the question “Does this image describe a place of banquet hall?" and the answer “Yes.", the negation argument is “This image does not describe a place of banquet hall." (2) Multiple-Choice Questions. For multiple-choice questions, negation arguments are produced by randomly selecting an incorrect option, and presenting it in a negated or contradicted form with the correct answer. Given the question and options, “What organism does this diagram portray? Options: A. ladybug B. wolf C. human D. plant", as well as the answer “D". The negation argument could be “The diagram portrays a ladybug." (option A). (3) Free-Form Questions. For free-form answers, negation arguments are crafted to contradict the provided factual answer. For instance, give the question “How many food items are shown in the bar graph? " and answer “14", the negation argument is “The number of food items shown in the bar graph is not 14."
Post-processing. To ensure consistency in evaluation, we employ a post-processing step to handle variations in the responses generated by MLLMs. Since the model outputs may not always align exactly with the expected format, we utilized Llama3-8B-Instruct and Qwen-14B-Chat to refine the responses. These models are tasked with interpreting the question, the expected answer, and the MLLM-generated response, normalizing the output to align with the desired format.
2.2 Results
Model Negation Dataset average
MME fu2023mme MMMU yue2023mmmu MMMUPro yue2024mmmu-MMMUPro AI2Diagram kembhavi2016diagram MathVista lu2023mathvista ChartQA masry2022chartqa PoPE li2023evaluating-Pope MMBench liu2025mmbench
Open-Source
before 79.70 31.77 12.37 59.52 33.10 51.72 86.24 68.79 52.90
[…]
before 86.02 50.37 27.05 78.37 60.00 75.32 87.94 82.26 68.42
[…]
before 70.01 42.98 12.14 67.81 34.90 60.72 88.96 79.44 57.12
[…]
before 91.70 60.34 33.64 83.42 67.10 80.60 87.84 79.81 73.06
[…]
Proprietary
before 82.43 57.39 31.98 71.99 48.70 68.37 80.27 78.84 65.00
[…]
before 68.45 60.47 43.12 78.30 59.00 72.28 63.60
[…]
before 84.46 60.34 51.56 70.63 60.30 78.84 67.69
[…]
Table 1: Performance comparison of selected Multimodal Large Language Models (MLLMs) across various benchmarks before (i.e., initial answers) and after the introduction of negation arguments. The performance drop is highlighted in red.
Performance Comparison. Table 1 lists the results across all evaluated benchmarks, comparing the accuracy of selected MLLMs before (i.e., initial answers) and after the introduction of negation arguments. Overall, all the MLLMs exhibit significant performance declines even though the accuracy of accuracy before negation is relatively high, with open-source models generally showing greater vulnerability compared to proprietary counterparts. The average accuracy drop ranges from 9.62% (LLaVa-1.6-7B in MMMUPro dataset) to 66.60 (Qwen2-VL-72B-Instruct in MMBench dataset), highlighting the widespread challenges MLLMs encounter when handling negation-based inputs in conversation.
Open-Source vs. Proprietary Models. Proprietary models generally showed better resilience. For instance, Gemini-1.5-flash has the lowest average accuracy drop 20.61%, maintaining 44.39% accuracy after negation, whereas open-source models like Qwen2-VL-72B-Instruct suffered dramatic declines, with a 46.15% drop and only manages to obtain the 26.91% accuracy on average. This highlights the need for improved adversarial training in open-source MLLMs to deal with the negation arguments in conversation. In addition, proprietary models like GPT-4o do not consistently outperform open-source models across all datasets. On tasks involving true/false or multiple-choice questions, such as those in MME and AI2Diagram, GPT-4o often fails to provide a definitive answer, frequently responding with “unknown." For instance, in MME, GPT-4o incorrectly answered nearly 150 out of 400 questions related to artwork, many of which were labeled as "unknown." Furthermore, GPT-4o struggles with datasets like MathVista and ChartQA, where large open-source models like Qwen2-VL-72B outperformed it.
Comparison between Larger and Smaller Models. Larger models like Qwen2-VL-72B-Instruct performed significantly better initially before negation, achieving an average accuracy of 73.06%, compared to smaller models like Qwen2-VL-7B-Instruct with 68.42%. However, larger models also exhibited more substantial drops after negation, with Qwen2-VL-72B-Instruct seeing a drastic 46.15% decline compared to Qwen2-VL-7B-Instruct’s 32.15% drop. This suggests that larger models may be more prone to overfitting or misalignment in adversarial scenarios like negation.
Qualitative Analysis. Figure 3 illustrates a series of examples showcasing how MLLMs respond to negation arguments in different datasets. In each example, the models initially provide correct responses. However, when users introduce negation arguments, many models revise their answers incorrectly. For instance, GPT-4o initially identifies the color representing New Zealand correctly but revises its answer after the negation argument. In addition, the models often generate detailed but fabricated explanations to justify their revised answers. For example, in the fourth example in MMMU dataset, Gemini-1.5-flash generates explanations following negation arguments lack grounding in the visual content, highlighting the models’ overconfidence in producing unverified reasoning.
Figure 3: Qualitative examples illustrating how various MLLMs respond to negation arguments when their initial answers are correct. Correct responses are highlighted in green, while incorrect responses are marked in red.
Model LLaVA-1.6-7B liu2024llavanext-llava1.6 LLaVA-NeXT-8B li2024llavanext-strong Qwen2-VL-7B-Instruct Qwen2VL-7B Gemini-1.5-flash team2024gemini1.5
before negation 31.77 42.98 50.37 57.39
[…]
Table 2: Performance comparison of selected Multimodal Large Language Models (MLLMs) under two distinct negation strategies from the MMMU dataset. The performance drop compared to before negation is indicated in red.
Effect of Negation Argument. To investigate the impact of different negation strategies, we conduct experiments using two distinct negation approaches in the MMMU dataset: (1) Direct Negation Using Options. This strategy refers to directly negating the initial answer by stating the correct option explicitly, such as “The correct answer is C." (2) Negation with Descriptive Content of Option represents that we negate the initial answer by providing a description of an incorrect option, such as “… is Explicit Themes (the content of option C)." As shown in Table 2, all the MLLMs show higher accuracy for the second strategy after negation. The result suggests that providing additional context can help models better retain reasoning integrity when challenged with negation arguments. Since descriptive content more effectively demonstrates the capabilities of MLLMs compared to direct options, we adopt the second strategy for all experiments. Additionally, we acknowledge that negation arguments can take numerous other forms, such as employing alternative phrasing methods. Exploring these variations will be our future work to comprehensively assess model vulnerabilities and enhance the robustness of MLLMs under diverse negation conditions.
3 Related Work
Multimodal Large Language Models. Multimodal Large Language Models (MLLMs) yin2023survey represent a significant evolution in artificial intelligence, integrating multiple data modalities—such as text, images, and audio—to enhance understanding and generation capabilities. The introduction of contrastive learning techniques in vision-language models such as CLIP radford2021learning enables cross-modal understanding through training on extensive datasets. Building upon these foundations, modern MLLMs like GPT-4o hurst2024gpt , Claude-3.5-Sonnet anthropic2024claude , and Gemini-1.5-flash team2024gemini1.5 have integrated vision encoders with large language models, achieving state-of-the-art performance in tasks including visual question answering and image reasoning. Open-source counterparts, such as LLaVA li2024llavanext-strong , Qwen2-VL Qwen2VL-7B and Deepseek liu2024deepseek , have democratized access to advanced multimodal technologies, fostering innovation within the research community. The typical training process for these models involves two stages: vision-language alignment pretraining, which maps visual features to the language model’s embedding space using large-scale image-text pairs, and visual instruction tuning, which fine-tunes the model to handle diverse visual instructions. Despite these advancements, this paper reveals that MLLMs are vulnerable to misleading negation arguments, even when their initial responses are correct. Addressing such weaknesses is essential for improving the reliability and robustness of MLLMs in practical applications.
Negation Understanding. Negation, defined as the contradiction or denial of something, is a fundamental aspect of language croft1991evolution ; pea1978development . In natural language processing (NLP), Pretrained Language Models like BERT have been shown to struggle with distinguishing affirmation and negation cloze questions kassner2019negated ; ettinger2020bert . To address this, methods like unlikelihood training have been introduced to improve BERT’s ability to understand negation using negated sentences from existing datasets hosseini2021understanding . This line of research has advanced further with studies like truong2023language , which show large language models (LLM), including GPT3 and InstructGPT, also exhibit significant limitations in handling negation. These models often fail to recognize the presence of negation, struggle to accurately interpret the lexical semantics of negation, and face challenges in maintaining logical consistency and reasoning effectively when confronted with negated contexts. Moreover, LLMs frequently struggle to maintain logical consistency when defending correct beliefs against invalid arguments, raising concerns about their alignment and depth of understanding wang2023can . Very recent work zhao2025aligning addresses the challenge of maintaining faithful integrity in large language models, which leverage bilateral confidence estimation and direct preference optimization to ensure LLMs remain consistent with correct statements while adapting to valid opposing arguments. Building on the insights of wang2023can ; zhao2025aligning , this paper extends the investigation into the impact of opposing arguments in MLLMs.
In multimodal learning, negation has received comparatively less attention than in NLP. Existing works have been explored to improve the ability to understand opposing arguments of CLIP-like vision-language models (VLMs) yuksekgonul2022and ; singh2024learn ; wang2023clipn ; alhamoud2025vision . For instance, very recent work alhamoud2025vision highlights the limitations of VLMs in understanding negation across tasks such as retrieval and multiple-choice questions with negation, while demonstrating that fine-tuning with negation-focused datasets can significantly improve performance. In contrast, this paper focuses on evaluating the negation understanding of MLLMs in conversational settings, particularly when the models’ original answers are correct but become misled by unfaithful negation arguments during conversation.
4 Conclusion
This paper has revealed critical vulnerabilities in MLLMs when exposed to negation arguments in conversation when the initial answers are correct. While these models demonstrate remarkable advancements in multimodal understanding and reasoning, their susceptibility to misleading prompts highlights fundamental weaknesses in their alignment and reasoning mechanisms. Through comprehensive evaluation across diverse benchmarks, we emphasize the importance of addressing these challenges to enhance the reliability and robustness of MLLMs. This paper also provides valuable insights for future research aimed at developing more trustworthy AI systems capable of maintaining logical consistency under adversarial conditions. We encourage further exploration of advanced training techniques and evaluation frameworks to mitigate the risks posed by such adversarial inputs, ultimately advancing the state of multimodal AI toward more reliable and trustworthy real-world applications.