### v2

# Calling a Spade a Heart: Gaslighting Multimodal Large Language Models via Negation

Abstract
Multimodal Large Language Models (MLLMs) have exhibited remarkable advancements in integrating different modalities, excelling in complex understanding and generation tasks. Despite their success, MLLMs remain vulnerable to conversational adversarial inputs, particularly negation arguments. This paper systematically evaluates state-of-the-art MLLMs across diverse benchmarks, revealing significant performance drops when negation arguments are introduced to initially correct responses. Notably, we introduce the first benchmark GaslightingBench, specifically designed to evaluate the vulnerability of MLLMs to negation arguments. GaslightingBench consists of multiple-choice questions curated from existing datasets, along with generated negation prompts across 20 diverse categories. Throughout extensive evaluation, we find that proprietary models such as Gemini-1.5-flash, GPT-4o and Claude-3.5-Sonnet demonstrate better resilience compared to open-source counterparts like Qwen2-VL and LLaVA. However, all evaluated MLLMs struggle to maintain logical consistency under negation arguments during conversation. Our findings provide critical insights for improving the robustness of MLLMs against negation inputs, contributing to the development of more reliable and trustworthy multimodal AI systems.
1 Introduction
Figure 1: Examples demonstrate that GPT-4o initially provides correct answers but incorrectly revises its responses when confronted with user-provided negation arguments. GPT-4o shows a tendency to accept misleading inputs, often generating hallucinated explanations to justify the revised answers, a behavior that can be described as a form of “gaslighting". Note: "Gaslighting is the manipulation of someone into questioning their own perception of reality." - Wikipedia.
Multimodal Large Language Models (MLLMs) have achieved remarkable progress in understanding and generating language grounded in multimodal contexts, such as visual and textual inputs [32, 19, 3, 10]. These models leverage cutting-edge advancements in Large Language Models (LLM) [36] and multimodal learning [25, 14, 17], enabling them to excel in diverse tasks, such as image understanding, visual question answering, and multimodal reasoning. Recent breakthroughs, such as the introduction of proprietary models GPT-4o [10] and Claude-3.5-Sonnet [2], have demonstrated exceptional performance, establishing new standards in multimodal understanding and reasoning. Additionally, open-source models like DeepSeek-V3 and DeepSeek-R1 [16, 7] are closing the performance gap with proprietary models, reflecting the rapid and competitive evolution of MLLMs. These advancements not only highlight the growing capabilities of LLMs, but also suggest that the development trajectory is edging closer to achieving Artificial General Intelligence (AGI).
Despite the impressive advancements, MLLMs exhibit significant vulnerabilities when navigating complex conversational challenges, particularly those involving adversarial negation. This issue becomes evident when models struggle to critically analyze and resist unfaithful arguments, resulting in erroneous reversals of their initially correct answers [29, 37]. As illustrated in Figure 1, GPT-4o initially provides correct answers for all examples. However, when presented with negation arguments by the user, GPT-4o often accepts these arguments, revising its answers incorrectly. Moreover, it even generates hallucinated explanations to justify these erroneous responses. Such susceptibility not only undermines their reliability but also exposes fundamental weaknesses in their reasoning and alignment mechanisms. These limitations are particularly concerning in high-stakes applications like healthcare diagnostics, autonomous decision-making, and content moderation, where the ability to maintain logical consistency and resist manipulation is crucial. Addressing these challenges requires a deeper understanding of how MLLMs process and align multimodal inputs, paving the way for more reliable and trustworthy models [9, 20].
Figure 2: Comparison of MLLMs’s performance before (i.e., initial answers) and after negation, reported as average accuracy across eight benchmarks-MME [6], MMMU [33], MMMUPro [34], MMBench [21], PoPE [15], ChartQA [23], AI2Diagram [12] and MathVista [22]). The results highlight the substantial accuracy drop across all models when negation is introduced. The access dates for GPT-4o and Claude-3.5-Sonnet are from 3-March-2025 to 7-March-2025, while results for Gemini-1.5-flash are based on access from 13-Jan-2025 to 18-Jan-2025. More detailed results are available in Table 1.
To comprehensively understand the limitations of MLLMs in handling negation, we conduct extensive evaluations across eight multimodal benchmarks and state-of-the-art models. The benchmarks cover diverse datasets, ranging from general multimodal datasets such as MMMU [33] and MMBench [21], chart dataset such as ChartQA [23], to math dataset MathVista [22], providing a robust framework for assessing the performance of MLLMs under negation arguments in conversation. The proprietary models, such as Gemini-1.5-flash, GPT-4o and Claude-3.5-Sonnet, alongside open-source counterparts like Qwen2-VL and LLaVA, allowed for a comparative analysis of reliability and accuracy. As shown in Figure 2, through the extensive evaluation, we highlight the widespread vulnerability of MLLMs to negation arguments, with varying degrees of impact observed across different datasets and MLLMs.
As existing benchmarks primarily evaluate factual accuracy and multimodal reasoning, they fail to systematically assess MLLMs’ susceptibility to negation arguments. To address this gap, we introduce GaslightingBench, the first multimodal benchmark designed to evaluate models’ ability to resist misleading negation arguments while maintaining logical consistency. We curate representative multiple-choice questions (MCQs) from established datasets and generated corresponding negation prompts, resulting in a collection spanning 20 categories and 1,287 samples.
Our observations reveal critical insights into the behavior of MLLMs. Notably, while larger models such as Qwen2-VL-72B-Instruct demonstrate higher accuracy in initial responses before negation, it also exhibit significant performance drops after negation, suggesting potential overfitting or alignment issues during training. On the contrary, proprietary models such as Gemini-1.5-flash, show better robustness compared to open-source models, but even they struggle to consistently defend correct answers against misleading negation arguments. Furthermore, different categories demonstrate varying levels of vulnerability to negation, with “Social Relation" suffering the most significant performance declines, whereas “Geography" experiences comparatively smaller yet still notable drops in our GaslightingBench. These findings emphasize the urgent need for enhanced training techniques and robust alignment mechanisms to improve the reliability and integrity of MLLMs in real-world applications.
2 Related Work
Multimodal Large Language Models. Multimodal Large Language Models (MLLMs) [32] represent a significant evolution in artificial intelligence, integrating multiple data modalities—such as text, images, and audio—to enhance understanding and generation capabilities. The introduction of contrastive learning techniques in vision-language models such as CLIP [25] enables cross-modal understanding through training on extensive datasets. Building upon these foundations, modern MLLMs like GPT-4o [10], Claude-3.5-Sonnet [2], and Gemini-1.5-flash [27] have integrated vision encoders with large language models, achieving state-of-the-art performance in tasks including visual question answering and image reasoning. Open-source counterparts, such as LLaVA [13] and Qwen2-VL [31], have democratized access to advanced multimodal technologies, fostering innovation within the research community. The typical training process for these models involves two stages: vision-language alignment pretraining, which maps visual features to the language model’s embedding space using large-scale image-text pairs, and visual instruction tuning, which fine-tunes the model to handle diverse visual instructions. Despite these advancements, this paper reveals that MLLMs are vulnerable to misleading negation arguments, even when their initial responses are correct. Addressing such weaknesses is essential for improving the reliability and robustness of MLLMs in practical applications.
Figure 3: Evaluation pipeline for assessing the robustness of Multimodal Large Language Models (MLLMs) to negation. The pipeline consists of three key stages: (1) Inputs and Initial Answers: MLLMs receive a variety of question formats as input, including Yes/No, Multiple-Choice, and Free-Form, and their initial answers are recorded. (2) Negation Generation: if the model’s initial response is correct, a negation argument is introduced to challenge its answer. Different negation strategies are applied based on the question type. (3) Post-Negation Evaluation: the model’s response after negation is analyzed to determine if it maintains consistency or is misled into revising its answer. Post-processing is applied to normalize responses for accurate comparison.
Negation Understanding. Negation, defined as the contradiction or denial of something, is a fundamental aspect of language [4, 24]. In natural language processing (NLP), Pretrained Language Models like BERT have been shown to struggle with distinguishing affirmation and negation cloze questions [11, 5]. To address this, methods like unlikelihood training have been introduced to improve BERT’s ability to understand negation using negated sentences from existing datasets [8]. This line of research has advanced further with studies like [28], which show large language models (LLM), including GPT3 and InstructGPT, also exhibit significant limitations in handling negation. These models often fail to recognize the presence of negation, struggle to accurately interpret the lexical semantics of negation, and face challenges in maintaining logical consistency and reasoning effectively when confronted with negated contexts. Moreover, LLMs frequently struggle to maintain logical consistency when defending correct beliefs against invalid arguments, raising concerns about their alignment and depth of understanding [29]. Very recent work [37] addresses the challenge of maintaining faithful integrity in large language models, which leverage bilateral confidence estimation and direct preference optimization to ensure LLMs remain consistent with correct statements while adapting to valid opposing arguments. Building on the insights of [29, 37], this paper extends the investigation into the impact of opposing arguments in MLLMs.
In multimodal learning, negation has received comparatively less attention than in NLP. Existing works have been explored to improve the ability to understand opposing arguments of CLIP-like vision-language models (VLMs) [35, 26, 30, 1]. For instance, very recent work [1] highlights the limitations of VLMs in understanding negation across tasks such as retrieval and multiple-choice questions with negation, while demonstrating that fine-tuning with negation-focused datasets can significantly improve performance. In contrast, this paper focuses on evaluating the negation understanding of MLLMs in conversational settings, particularly when the models’ original answers are correct but become misled by unfaithful negation arguments during conversation.
3 Evaluating Multimodal Large Language Model via Negation
3.1 Experimental Setup
Figure 4: The category distribution of GaslightingBench with 20 categories and 1,287 samples. Each category is carefully curated from existing datasets to ensure balanced representation and broad coverage, providing a comprehensive evaluation dataset for assessing MLLMs’ vulnerabilities to negation arguments.
Figure 5: Examples from different categories in the GaslightingBench. The highlighted green option is the correct answer, while a randomly selected incorrect option is used to generate the negation argument.
Multimodal Large Language Models. To establish a comprehensive evaluation of the negation challenges posed to Multimodal Large Language Models (MLLMs), we assess a range of state-of-the-art models representing both proprietary and open-source systems. These models were chosen to reflect a diverse set of training approaches and capabilities in the multimodal AI landscape.
Proprietary Models: Our evaluation included leading proprietary models such as GPT-4o [10], Claude-3.5-Sonnet [2], and Gemini-1.5-flash [27]. These models represent the most powerful multimodal AI capabilities.
Open-source Models: In addition to proprietary models, we included open-source models such as LLaVA-1.6-7B [18], Qwen2-VL-7B-Instruct [31], LLaVA-NeXT-8B [13] and Qwen2-VL-72B-Instruct [31]. These open-source models showcase the rapid progress in publicly available multimodal AI systems.
Benchmarks. To comprehensively assess the robustness of MLLMs in handling negation, we employ a diverse set of benchmark datasets, including general multimodal datasets MME [6], MMMU [33], MMMUPro [34], MMBench [21] and PoPE [15], Chart dataset ChartQA [23], diagram dataset AI2Diagram [12] as well as Math dataset MathVista [22].
3.2 Evaluation Pipeline
As shown in Figure 3, we introduce a structured evaluation pipeline to systematically assess MLLMs’ vulnerability to negation by measuring their performance before and after exposure to negation arguments. The process consists of three key steps: First, questions and corresponding options (if applicable) from the original datasets are provided to the MLLMs, and their initial responses are recorded. Second, if the model’s initial answer aligns with the correct answer from the dataset, we introduce a negation argument challenging the validity of the original answer. The model’s response to the negation argument is recorded to assess whether it revises its original correct answer. If the model’s initial answer is incorrect, no negation argument is introduced, and the response remains unchanged. Third, we compute the accuracy of the model both before and after the introduction of negation arguments to quantify the drop in performance caused by adversarial negation.
Negation Generation. As illustrated in Figure 3, we generate negation arguments tailored to different question formats using Llama3-8B-Instruct: (1) Yes/No Questions. Negation arguments are generated by rewriting the original question into an opposite statement based on the answer. If the answer is “Yes.", the question is rewritten as a negative statement and vice versa. For example, given the question “Is this photo of Grimspound?" and the answer “Yes.", the negation argument is “This it not a photo of Grimspound." (2) Multiple-Choice Questions. For multiple-choice questions, negation arguments are produced by randomly selecting an incorrect option, and presenting it in a negated or contradicted form with the correct answer. Given the question and options, “Which direction is the cat facing? Options: A. right B. upward C. left D. the camera", as well as the answer “D". The negation argument could be “The cat is facing right." (option A). (3) Free-Form Questions. For free-form answers, negation arguments are crafted to contradict the provided factual answer. For instance, give the question “What is the maximum value of y? " and answer “5", the negation argument is “The maximum value of y is not 5."
Model Negation Dataset average
MME [6] MMMU [33] MMMUPro [34] AI2Diagram [12] MathVista [22] ChartQA [23] PoPE [15] MMBench [21]
Open-Source
before 79.70 31.77 12.37 60.40 33.10 51.72 86.24 68.79 53.01
[…]
before 86.02 50.37 27.05 79.83 60.00 75.32 87.94 82.26 68.42
[…]
before 70.01 42.98 12.14 68.56 34.90 60.72 88.96 79.44 57.21
[…]
before 91.70 60.34 33.64 85.30 67.10 80.60 87.84 79.81 73.29
[…]
Proprietary
before 82.43 57.39 31.98 73.52 48.70 68.37 80.27 78.84 65.19
[…]
before 69.17 62.07 36.42 75.84 54.60 72.68 85.46 77.69 66.74
[…]
before 86.60 67.73 37.40 72.51 57.50 77.40 81.71 83.07 70.49
[…]
Table 1: Performance comparison of Multimodal Large Language Models (MLLMs) across various benchmarks before (i.e., initial answers) and after the introduction of negation arguments. The performance drop is highlighted in red.
Model LLaVA-1.6-7B [18] Qwen2-VL-7B-Instruct [31] LLaVA-NeXT-8B [13] Qwen2-VL-72B-Instruct [31] Gemini-1.5-flash [27] GPT-4o [10] Claude-3.5-Sonnet [2]
before negation 59.13 76.85 71.33 77.08 74.13 69.15 78.17
[…]
Table 2: Results of MLLMs in our GaslightingBench, comparing each model’s performance before (i.e., initial answers) and after negation. The performance drop is highlighted in red.
Model LLaVA-1.6-7B [18] LLaVA-NeXT-8B [13] Qwen2-VL-7B-Instruct [31] Gemini-1.5-flash [27]
before negation 31.77 42.98 50.37 57.39
[…]
Table 3: Performance comparison of Multimodal Large Language Models (MLLMs) under two distinct negation strategies from the MMMU dataset. The performance drop compared to before negation is indicated in red.
Post-processing. To ensure consistency in evaluation, we employ a post-processing step to handle variations in the responses generated by MLLMs. Since the model outputs may not always align exactly with the expected format, we utilized Llama3-8B-Instruct and Qwen-14B-Chat to refine the responses. These models are tasked with interpreting the question, the expected answer, and the MLLM-generated response, normalizing the output to align with the desired format.
3.3 Gaslighting Benchmark Construction
To comprehensively evaluate the impact of negation on MLLMs, we introduce the first gaslighting benchmark, GaslightingBench, designed to ensure broad coverage, category balance, and the ability to expose MLLMs’ vulnerabilities to negation arguments. GaslightingBench exclusively uses Multiple-Choice Questions (MCQ) to facilitate structured evaluation, as MLLMs generally provide more consistent responses in MCQs than in open-ended formats, while MCQs remain more complex than binary Yes/No questions. Our benchmark is constructed through a two-step process. We first select representative questions and images. MCQs are extracted from existing datasets following three key principles: (1) We carefully review the categories across all the existing benchmarks and select a balanced representation of general categories. (2) To maintain a sufficient number of samples per category, we manually merge semantically similar categories. For example, in MMBench,“Attibute_comparison" and “Attribute_recognition" are combined under “Attribute" to ensure broader coverage. (3) For categories present in multiple datasets, we incorporate samples from different sources to ensure representativeness. For instance, our “Math" category consists of samples from both AI4MATH and MMMU-Math. Second, we employ the same approach detailed in section 3.2 to generate negation arguments for all the selected questions. Ultimately, as shown in Figure 4, GaslightingBench comprises 20 categories and 1,287 samples, covering a wide range of topics, such as Business, Medicine, Image Emotion, and Counting. Figure 5 shows a few examples from different categories in the GaslightingBench.
4 Result Analysis
Figure 6: Accuracy drop across different categories in GaslightingBench.
4.1 Performance Comparison
Table 1 lists the results across all evaluated benchmarks, comparing the accuracy of selected MLLMs before (i.e., initial answers) and after the introduction of negation arguments. Overall, all the MLLMs exhibit significant performance declines even though the accuracy before negation is relatively high, with open-source models generally showing greater vulnerability compared to proprietary counterparts. The average accuracy drop ranges from 9.62% (LLaVa-1.6-7B in MMMUPro dataset) to 66.60% (Qwen2-VL-72B-Instruct in MMBench dataset), highlighting the widespread challenges MLLMs encounter when handling negation-based inputs in conversation.
Open-Source vs. Proprietary Models. Proprietary models generally showed better resilience. For instance, Gemini-1.5-flash has the lowest average accuracy drop 20.53%, maintaining 44.66% accuracy after negation, whereas open-source models like Qwen2-VL-72B-Instruct suffered dramatic declines, with a 46.02% drop and only manages to obtain 27.27% accuracy on average. This highlights the need for improved adversarial training in open-source MLLMs to deal with the negation arguments in conversation. In addition, proprietary models like GPT-4o do not consistently outperform open-source models across all datasets. On tasks involving true/false or multiple-choice questions, such as those in MME and AI2Diagram, GPT-4o often fails to provide a definitive answer, frequently responding with “unknown." For instance, in MME, GPT-4o incorrectly answered nearly 150 out of 400 questions related to artwork, many of which were labeled as "unknown." Furthermore, GPT-4o struggles with datasets like MathVista and ChartQA, where large open-source models like Qwen2-VL-72B outperformed it.
Comparison between Larger and Smaller Models. Larger models like Qwen2-VL-72B-Instruct performed significantly better initially before negation, achieving an average accuracy of 73.29%, compared to smaller models like Qwen2-VL-7B-Instruct with 68.42%. However, larger models also exhibited more substantial drops after negation, with Qwen2-VL-72B-Instruct seeing a drastic 46.02% decline compared to Qwen2-VL-7B-Instruct’s 32.17% drop. This suggests that larger models may be more prone to overfitting or misalignment in adversarial scenarios like negation.
Figure 7: Qualitative examples illustrating how various MLLMs respond to negation arguments when their initial answers are correct. Correct responses are highlighted in green, while incorrect responses are marked in red.
Result in GaslightingBench. Table 2 presents the evaluation results of MLLMs on GaslightingBench, highlighting their accuracy before and after the introduction of negation arguments. Similarly, across all models, significant performance degradation is observed, reaffirming the susceptibility of MLLMs to negation. As shown in Figure 6, different categories show varying degrees of susceptibility to negation arguments. Notably, “Social Relation" and “Image Emotion" experience the most severe drops, with accuracy declining by over 50%, indicating their high sensitivity to negation arguments. In contrast, “Geography" and “Chemistry" demonstrate comparatively lower declines but still suffer more than a 25% accuracy reduction, emphasizing that no category remains unaffected. Similar observations also appear in other existing datasets, for instance, MMMU dataset, which contains 30 subjects such as “accounting" and “Art”, the largest performance drop is observed in “Art Theory", while “Music" experiences the smallest decline. Similarly, in the MME dataset, “Counting" shows the most significant drop, whereas “Numerical Calculation" is the least affected. These results highlight the uneven but widespread impact of negation arguments across different domains.
Qualitative Analysis. Figure 7 illustrates examples showcasing how MLLMs respond to negation arguments in different datasets. In each example, the models initially provide correct responses. However, when users introduce negation arguments, many models revise their answers incorrectly. For instance, in the first example in ChartQA dataset, GPT-4o initially identifies the color representing New Zealand correctly but revises its answer after the negation argument. In addition, the models often generate detailed but fabricated explanations to justify their revised answers. For example, in the second example in MMMU dataset, LLaVA-1.6-7B generates explanations following negation arguments lack grounding in the visual content, highlighting the models’ overconfidence in producing unverified reasoning.
Figure 8: Example shows that different options in negation arguments (e.g., “ten," “one," and “eight") affect MLLMs’ ability to maintain their initial correct response.
4.2 Discussion: Effect of Negation Argument
To investigate the impact of different negation strategies, we conduct experiments using two distinct negation approaches in the MMMU dataset: (1) Direct Negation Using Options. This strategy refers to directly negating the initial answer by stating the correct option explicitly, such as “The correct answer is C." (2) Negation with Descriptive Content of Option represents that we negate the initial answer by providing a description of an incorrect option, such as “… is Explicit Themes (the content of option C)." As shown in Table 3, all the MLLMs show higher accuracy for the second strategy after negation. The result suggests that providing additional context can help models better retain reasoning integrity when challenged with negation arguments. Since descriptive content more effectively demonstrates the capabilities of MLLMs compared to direct options, we adopt the second strategy for all experiments. Furthermore, as shown in Figure 8, the choice of negation arguments significantly affects MLLMs’ performance. For instance, when the negation argument includes “ten" or “eight," only two models maintain the correct response of “four," whereas using “one" as the negation argument allows six models to retain the correct answer. This highlights the varying sensitivity of MLLMs to different negation formulations. Additionally, we acknowledge that negation arguments can take numerous other forms, such as employing alternative phrasing methods. Exploring these variations will be our future work to comprehensively assess model vulnerabilities and enhance the robustness of MLLMs under diverse negation conditions.
5 Conclusion
This paper has revealed critical vulnerabilities in MLLMs when exposed to negation arguments in conversation when the initial answers are correct. While these models demonstrate remarkable advancements in multimodal understanding and reasoning, their susceptibility to misleading prompts highlights fundamental weaknesses in their alignment and reasoning mechanisms. Through comprehensive evaluation across diverse benchmarks, we emphasize the importance of addressing these challenges to enhance the reliability and robustness of MLLMs. This paper also provides valuable insights for future research aimed at developing more trustworthy AI systems capable of maintaining logical consistency under adversarial negation conditions. We encourage further exploration of advanced training techniques and evaluation frameworks to mitigate the risks posed by such adversarial inputs, ultimately advancing the state of multimodal AI toward more reliable and trustworthy real-world applications.